2kg of gas at a pressure of 1.5 bar.occupies a volume of 2.5m cube.if this gas compresses isothermally to 1/3 times the initial volume.find initial.final temperature,work done,heat transfer.

Grade:12th pass

1 Answers

Vikas TU
7 years ago
As the pocess is isothermal chanege in Internal energy would be zero and hence temperature remains costant that is T.
From ideal gas eqn.
PV =nRT
1.5 * 2.5*1000 = 2000/32 * 0.0821 * T
we get T = 730.8 K
 
W = 2.303*nRTlog(Vf/Vi)
    = 2.303*(62.5)*0.0821*730.8*log(3)
    = 4120.44 KJ
 
Heat Transfer = W = 4120.44 KJ.